OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given = 6sqre root two cm

OpenStudy (anonymous):

How do I solve?

OpenStudy (anonymous):

Use the hypotenuse formula to solve this.

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\]?

OpenStudy (anonymous):

A^2 + B^2 = C^2. A= leg B= leg C= Hypothenuse Basic rule derived from: The Pythagoreom Theorem Also 3+4=5. Its a set to help you out.

OpenStudy (anonymous):

so would it be 36+36=c^2?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

my online school doesn't agree with the answer I got.

OpenStudy (anonymous):

Then they have an another solution to do this

OpenStudy (anonymous):

 For any isosceles right triangle, these will be the measures of the sides. You can prove it with the Pythagorean theorem